# Adapted optimal transport between Gaussian processes in discrete time

### v1

@@ 4. Dynamic programming principle
  • Case 3: $\alpha_{t+1}=(L^{\top}M)_{t+1,t+1}=0$. The value of the integral in (4.9) is independent of the choice of $\pi_{t+1}$.
By Proposition 4.2, $\pi_{t+1}=\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is optimal for (4.9). We have, by the induction hypothesis, tower property and Proposition 4.2 again,
$\begin{split}&V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int\int\|x-y\|^{2}\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{\mathbf{t}_{+}^{\prime}}(\mathrm{d}x_{\mathbf{t}_{+}^{\prime}},\mathrm{d}y_{\mathbf{t}_{+}^{\prime}}|x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\mathrm{d}x_{t+1},\mathrm{d}y_{t+1}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int\|x-y\|^{2}\pi_{\mathrm{AW}}^{\mu,\nu}(\mathrm{d}x_{\mathbf{t}^{\prime}},\mathrm{d}y_{\mathbf{t}^{\prime}}|x_{\mathbf{t}},y_{\mathbf{t}}),\end{split}$
which is the desired expression. In particular, we may express (4.6) as
$V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})=\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+\mathcal{AW}_{2}^{2}(\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}})).$
@@
Let $\mu=\mathcal{N}(a,A)$ and $\nu=\mathcal{N}(b,B)$ be non-degenerate Gaussian distributions on $\mathbb{R}^{N}$. For $\pi\in\Pi_{bc}(\mu,\nu)$, decompose it as a product of conditional distributions as in (2.5). Then $\pi$ is optimal for $\mathcal{AW}_{2}(\mu,\nu)$ if and only if for almost all $x_{1:(t-1)}$ and $y_{1:(t-1)}$, $\pi_{t}(\cdot|x_{1:(t-1)},y_{1:(t-1)})$ is the comonotonic (resp. counter-monotonic) coupling between $\mu_{t}(\cdot|x_{1:(t-1)})$ and $\nu_{t}(\cdot|x_{1:(t-1)})$ when $(L^{T}M)_{tt}>0$ (resp. $(L^{T}M)_{tt}<0$). [^3] We note again that $(L^{\top}M)_{NN}>0$. The optimal coupling is unique if and only if $(L^{\top}M)_{tt}\neq 0$ for all $t$. In this case it is deterministic and is given by
@@
where $P$ is the diagonal matrix with $P_{st}=\delta_{st}\rho_{t}$ satisfying (3.7). In all cases (any version of) $\pi_{\mathrm{AW}}^{\mu,\nu}$ is optimal.
@@
Our proof of Theorem 1.1 depends heavily on the Gaussian assumption and the resulting linear-quadratic structure of the value function (note that $V_{N}(x,y)=c(x,y)=\|x-y\|^{2}$ is the quadratic cost). Also, the signs $\rho_{t}$ are constants which do not depend on the pasts of $X$ and $Y$ (otherwise the DPP would be messier). While it may be possible to extend our techniques to other distributions such as elliptical distributions [13], in this paper we focus on the Gaussian case for concreteness and tractability. [^4] We thank Ludger Rüschendorf for his suggestion to consider elliptical distributions. Here we note that although the conditional distribution of an elliptical distribution remains elliptical and its mean vector and scale matrix are analogous to those in (4.4), the conditional covariance matrix depends on the conditioning value unless the distribution is Gaussian (see [22, page 36]). So Theorem 1.1 and Corollary 4.6 do not extend trivially. On the other hand, we note that there are no known analytical formulas for the $\mathcal{W}_{2}$-optimal transport map between generic elliptical distributions unless they share a common generator (in which case the optimal map has the form (2.4) [15]).
@@ 5. Adapted Wasserstein geometry of Gaussian distributions
In this section we discuss briefly the adapted Wasserstein geometry of Gaussian distributions, and leave a deeper study, in the spirits of [9], [11] and [26], to future research.
It is known that $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{AW}_{2})$, unlike $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{W}_{2})$, is not a complete metric space when $N\geq 2$. We observe that this is the case even if we restrict to Gaussian distributions. See [5, Section 4] for the completion of $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{AW}_{2})$ which involves the concept of nested distributions.
@@
Let $\mathscr{G}(N)=\{\mathcal{N}(a,A):a\in\mathbb{R}^{N},A\in\mathscr{S}_{+}(N)\}$ be the space of (possibly degenerate) Gaussian distributions on $\mathbb{R}^{N}$. Then $(\mathscr{G}(N),\mathcal{AW}_{2})$ is not complete when $N\geq 2$.
@@
$\mathcal{AW}_{2}^{2}(\mu_{n}(\theta),\mu_{m}(\theta))\leq d_{\mathrm{KR}}^{2}(A_{n}(\theta),A_{m}(\theta))=\|L_{n}(\theta)-L_{m}(\theta)\|_{\mathrm{F}}^{2}=2\left(\frac{1}{n}-\frac{1}{m}\right)^{2},$
@@
Next we consider geodesics. It is well known that the geodesic $(\mu_{t})_{0\leq t\leq 1}$ between non-degenerate $\mu_{0}=\mathcal{N}(a_{0},A_{0})$ and $\mu_{1}=\mathcal{N}(a_{1},A_{1})$ in $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{W}_{2})$ is given by McCann’s displacement interpolation [21]: $\mu_{t}=\mathcal{N}(a_{t},A_{t})$, where $(1-t)a_{0}+ta_{1}$,
@@
Let $A_{0},A_{1}\in\mathscr{S}_{++}(N)$. Define $A_{t}$, $0\leq t\leq 1$, by (5.2) where $T=L_{1}PL_{0}^{-1}$ as in (4.10), $P$ satisfies (3.7) and $\rho_{t}\in\{-1,+1\}$ when $(L^{\top}M)_{tt}=0$. Then $\mu_{t}=\mathcal{N}(a_{t},A_{t})$, where $a_{t}=(1-t)a_{0}+ta_{1}$, is a geodesic in $(\mathscr{G}(N),\mathcal{AW}_{2})$ between $\mu_{0}$ and $\mu_{1}$.
@@
On some probability space, let $X_{0}\sim\mathcal{N}(a_{1},A_{0})$. For $0\leq t\leq 1$, let $X_{t}=a_{t}+T_{t}(X_{0}-a_{0})$. Since $T_{t}$ is lower triangular and invertible, $(X_{s},X_{t})$ is a bicausal coupling of $(\mu_{s},\mu_{t})$ for $0\leq s,t\leq 1$. From Corollary 4.4, $(X_{0},X_{1})$ is $\mathcal{AW}_{2}$-optimal for $(\mu_{0},\mu_{1})$. From (5.2) we have, for $0\leq s,t\leq 1$,
@@
Since $s$ and $t$ are arbitrary, the triangle inequality implies that equality holds everywhere. ∎
We note that in Proposition 5.2 the $\mathcal{AW}_{2}$-geodesic is not necessarily unique when $L_{0}^{\top}L_{1}$ contains a zero entry. In fact, when $L_{0}^{\top}L_{1}$ contains a zero entry the interpolant $A_{t}$ may become degenerate, see Example 6.1.
It is well-known that the Bures-Wasserstein distance $d_{\mathrm{BW}}$ is a Riemannian distance on $\mathscr{S}_{++}(N)$ [26]; the Riemannian metric is Otto’s metric [23] restricted to the space of (centered) Gaussian distributions. From Corollary 3.2, the Knothe-Rosenblatt distance $d_{\mathrm{KR}}$ is Euclidean when expressed in terms of the Cholesky matrix. From (1.2) and (3.4), $d_{\mathrm{ABW}}$ and $d_{\mathrm{KR}}$ are related by
@@
From Corollary 4.6 (or directly from (5.3)), we have $d_{\mathrm{ABW}}(A,B)=d_{\mathrm{KR}}(A,B)$ when $A$ and $B$ are sufficiently close. Hence a Riemannian metric, if it existed, must be the Euclidean metric on $\mathscr{L}_{++}(N)$ under the Cholesky coordinate system ($A=LL^{\top}\mapsto L$), but $d_{\mathrm{ABW}}\not\equiv d_{\mathrm{KR}}$. ∎
@@ 6. Examples
In the following examples we let $\mu=\mathcal{N}(0,A)$ and $\nu=\mathcal{N}(0,B)$.
@@
[…]
@@
Figure 2. Three interpolations between $A_{0}=A$ and $A_{1}=B$ as in (6.1). The thin ellipses visualize McCann’s displacement interpolation. The thick dashed ones show the Knothe-Rosenblatt interpolation ($T=L_{1}L_{0}^{-1}$). The thick solid ones show the interpolation with $T=L_{1}PL_{0}^{-1}$; it is degenerate when $t=\frac{1}{2}$.
Next we show an example where the optimal $\mathcal{AW}_{2}$-coupling is not unique.
@@
[…]
@@
$X_{2}=X_{1}+\epsilon_{2},\quad Y_{1}=-Y_{1}+\epsilon_{2},$
where $\epsilon_{2}\sim\mathcal{N}(0,1)$ is the common noise at time $2$. Since

### v4

@@ 4. Dynamic programming principle
  • Case 3: $\alpha_{t+1}=(L^{\top}M)_{t+1,t+1}=0$. The value of the integral in (4.10) is independent of the choice of $\pi_{t+1}$. Hence any $\pi_{t+1}$ is optimal.
By Proposition 4.2, $\pi_{t+1}=\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\cdot|x_{\mathbf{t}},y_{\mathbf{t}})$ is optimal for (4.10). Using the induction hypothesis, tower property and Proposition 4.2, we have
$\begin{split}&V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int V_{t+1}(x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\mathrm{d}x_{t+1}\mathrm{d}y_{t+1}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int\left(\int\|x-y\|^{2}\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{\mathbf{t}_{+}^{\prime}}(\mathrm{d}x_{\mathbf{t}_{+}^{\prime}}\mathrm{d}y_{\mathbf{t}_{+}^{\prime}}|x_{\mathbf{t}_{+}},y_{\mathbf{t}_{+}})\right)\left(\pi_{\mathrm{AW}}^{\mu,\nu}\right)_{t+1}(\mathrm{d}x_{t+1}\mathrm{d}y_{t+1}|x_{\mathbf{t}},y_{\mathbf{t}})\\ &=\int\|x-y\|^{2}\pi_{\mathrm{AW}}^{\mu,\nu}(\mathrm{d}x_{\mathbf{t}^{\prime}}\mathrm{d}y_{\mathbf{t}^{\prime}}|x_{\mathbf{t}},y_{\mathbf{t}}),\end{split}$
which is the desired expression and completes the induction. In particular, we may express the value function (4.6) as $V_{t}(x_{\mathbf{t}},y_{\mathbf{t}})=\|x_{\mathbf{t}}-y_{\mathbf{t}}\|^{2}+\mathcal{AW}_{2}^{2}(\mu_{\mathbf{t}^{\prime}}(\cdot|x_{\mathbf{t}}),\nu_{\mathbf{t}^{\prime}}(\cdot|y_{\mathbf{t}}))$, with the convention that the last term is 00 when $t=N$. ∎
@@
Let $\mu=\mathcal{N}(a,A=LL^{\top})$ and $\nu=\mathcal{N}(b,B=MM^{\top})$ be non-degenerate Gaussian distributions on $\mathbb{R}^{N}$. For $\pi\in\Pi_{bc}(\mu,\nu)$, decompose it as a product of regular conditional distributions as in (2.5). Then $\pi$ is optimal for $\mathcal{AW}_{2}(\mu,\nu)$ if and only if for $\pi$-almost all $x_{1:(t-1)}$ and $y_{1:(t-1)}$, $\pi_{t}(\cdot|x_{1:(t-1)},y_{1:(t-1)})$ is the comonotonic (resp. counter-monotonic) coupling between $\mu_{t}(\cdot|x_{1:(t-1)})$ and $\nu_{t}(\cdot|x_{1:(t-1)})$ when $(L^{T}M)_{tt}>0$ (resp. $(L^{T}M)_{tt}<0$). The optimal coupling is unique if and only if $(L^{\top}M)_{tt}\neq 0$ for all $t$. In this case it is deterministic and is given by
@@
where $P$ is the diagonal matrix with $P_{st}=\delta_{st}\rho_{t}$ satisfying (3.6). In all cases (any version of) $\pi_{\mathrm{AW}}^{\mu,\nu}$, under which $(X,Y)$ is jointly Gaussian (possibly degenerate), is optimal.
@@
  • (i)
    The proof of Theorem 1.1 can be extended easily to weighted square losses of the form $c(x_{1:N},y_{1:N})=\sum_{t=1}^{N}w_{t}(x_{t}-y_{t})^{2}$ where $w_{t}>0$. Let $W=(\delta_{st}w_{t})$ be diagonal. Then the optimal coefficient $\rho_{t}$ becomes the sign of $(L^{\top}WM)_{tt}$, and the value of the bicausal OT problem $\inf_{\pi\in\Pi_{bc}(\mu,\nu)}\int_{\mathbb{R}^{N}\times\mathbb{R}^{N}}c(x,y)\pi(\mathrm{d}x\mathrm{d}y)$ becomes
    $(a-b)^{\top}W(a-b)+\operatorname{tr}(L^{\top}WL)+\operatorname{tr}(M^{\top}WM)-2\|\operatorname{diag}(L^{\top}WM)\|_{1}.$
  • (ii)
    Our techniques depend heavily on the Gaussian assumption and the linear-quadratic structure of the value function. As shown recently in [1], this structure extends to the multivariate setting with entropic regularization. While it may be possible to derive the optimal couplings for other distributions such as elliptical distributions [17], in this paper we focus on the Gaussian case for concreteness and tractability. [^5] We thank Ludger Rüschendorf for his suggestion to consider elliptical distributions. Here we note that although the conditional distribution of an elliptical distribution remains elliptical and the conditional mean vector and scale matrix are analogous to those in (4.3), the conditional covariance matrix depends on the conditioning value unless the distribution is Gaussian (see [24, page 36]). So Theorem 1.1 and Corollary 4.6 do not extend trivially.
We end this section with a necessary and sufficient condition for the Knothe-Rosenblatt coupling between non-degenerate Gaussian distributions to be $\mathcal{AW}_{2}$-optimal.
@@
Remark 4.7.
It was proved in [9, Proposition 3.4] that whenever $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{N})$ satisfy a stochastic co-monotonicity condition the Knothe-Rosenblatt coupling is $\mathcal{AW}_{2}$-optimal. Even for Gaussian (and Markovian) distributions, our necessary and sufficient condition in Corollary 4.6 is not equivalent to the stochastic co-monotonicity condition. We illustrate this in Example 6.3.
@@ 5. Adapted Wasserstein geometry of Gaussian distributions
We discuss briefly the adapted Wasserstein geometry of univariate Gaussian processes, and leave a deeper study, in the spirits of [7, 11, 14, 29], to future research. Let $\mathscr{G}(N)=\{\mathcal{N}(a,A):a\in\mathbb{R}^{N},A\in\mathscr{S}_{+}(N)\}$ be the space of (possibly degenerate) Gaussian distributions on $\mathbb{R}^{N}$.
It is known that $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{AW}_{2})$, unlike $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{W}_{2})$, is not a complete metric space when $N\geq 2$. Its completion can be identified with the space of nested distributions [7, Section 4] or the space of filtered processes [11]. Here, we observe that the subspace $(\mathscr{G}(N),\mathcal{AW}_{2})$ of Gaussian distributions is still incomplete.
@@
$(\mathscr{G}(N),\mathcal{AW}_{2})$ is not complete when $N\geq 2$.
@@
$\mathcal{AW}_{2}^{2}(\mu_{n}(\theta),\mu_{m}(\theta))\leq\mathcal{KR}_{2}^{2}(\mu_{n}(\theta),\mu_{m}(\theta))=\|L_{n}(\theta)-L_{m}(\theta)\|_{\mathrm{F}}^{2}=2\left(\frac{1}{n}-\frac{1}{m}\right)^{2},$
@@
We explain the probabilistic intuition of the construction in the proof of Proposition 5.1. Write $X=L(\theta)\epsilon$, so that $X_{1}=0$ and $X_{2}=(\cos\theta)\epsilon_{1}+(\sin\theta)\epsilon_{2}$. If we use the natural filtration $\mathcal{F}^{X}$ of $X$, $X_{1}=0$ contains no information about $X_{2}$. However, only $\epsilon_{2}$ is independent of $\mathcal{F}_{1}^{\epsilon}$ and the “angle” $\theta$ controls the relative exposure. Although $\mathcal{F}^{X}=\mathcal{F}^{\epsilon}$ when $X=L_{n}(\theta)\epsilon$, the equality between the two filtrations is not preserved in the limit.
Next we consider geodesics. [^6] A curve $(\gamma_{t})_{0\leq t\leq 1}$ is said to be a (constant-speed) geodesic on a metric space $(S,d)$ if $d(\gamma_{s},\gamma_{t})=|s-t|d(\gamma_{0},\gamma_{1})$ for all $s,t\in[0,1]$. It is well known that the unique geodesic $(\mu_{t})_{0\leq t\leq 1}$ between non-degenerate $\mu_{0}=\mathcal{N}(a_{0},A_{0})$ and $\mu_{1}=\mathcal{N}(a_{1},A_{1})$ in $(\mathcal{P}_{2}(\mathbb{R}^{N}),\mathcal{W}_{2})$ is given by McCann’s displacement interpolation [23]: $\mu_{t}=\mathcal{N}(a_{t},A_{t})$, where $a_{t}=(1-t)a_{0}+ta_{1}$,
@@
Let $A_{0},A_{1}\in\mathscr{S}_{++}(N)$. For $0\leq t\leq 1$, Define $A_{t}$ by (5.2) where $T=L_{1}PL_{0}^{-1}$ as in (4.11), $P$ satisfies (3.6) and $\rho_{t}\in\{-1,+1\}$ when $(L_{0}^{\top}L_{1})_{tt}=0$. Then $\mu_{t}=\mathcal{N}(a_{t},A_{t})$, where $a_{t}=(1-t)a_{0}+ta_{1}$, is a geodesic in $(\mathscr{G}(N),\mathcal{AW}_{2})$.
@@
On some probability space, let $X_{0}\sim\mathcal{N}(a_{0},A_{0})$. For $0\leq t\leq 1$, let $X_{t}=a_{t}+T_{t}(X_{0}-a_{0})$. Since $T_{t}$ is lower triangular and invertible, $(X_{s},X_{t})$ is a bicausal coupling of $(\mu_{s},\mu_{t})$ for $0\leq s,t\leq 1$. From Corollary 4.4, $(X_{0},X_{1})$ is $\mathcal{AW}_{2}$-optimal for $(\mu_{0},\mu_{1})$. From (5.2) we have, for $0\leq s,t\leq 1$,
@@
Since $s$ and $t$ are arbitrary, the triangle inequality implies that equality holds. ∎
We note that in Proposition 5.2 the $\mathcal{AW}_{2}$-geodesic is not necessarily unique when $L_{0}^{\top}L_{1}$ contains a zero entry. In fact, when $L_{0}^{\top}L_{1}$ contains a negative entry the interpolant $A_{t}$ may become degenerate, see Example 6.2.
It is well-known that the Bures-Wasserstein distance $d_{\mathrm{BW}}$ is a Riemannian distance on $\mathscr{S}_{++}(N)$ [29]; the Riemannian metric is Otto’s metric [25] restricted to the space of centered Gaussian distributions. From Corollary 3.2, the Knothe-Rosenblatt distance $d_{\mathrm{KR}}$ is Euclidean when expressed in terms of the Cholesky matrix. From (1.2) and (3.4), $d_{\mathrm{ABW}}$ and $d_{\mathrm{KR}}$ are related by
@@
We argue by contradiction. Suppose $d_{\mathrm{ABW}}$ is distance induced by a Riemannian metric on $\mathscr{S}_{++}(N)$. The metric tensor at $A\in\mathscr{S}_{++}(N)$ can be recovered by the distances $d_{\mathrm{ABW}}(A,B)$ where $B$ ranges over a neighborhood of $A$ [31, Example 9]. From Corollary 4.6, we have $d_{\mathrm{ABW}}(A,B)=d_{\mathrm{KR}}(A,B)$ when $A$ and $B$ are sufficiently close. This implies that the metric tensor must be that of $d_{\mathrm{KR}}$ (which is Euclidean under the Cholesky coordinate system $A\mapsto L$). This is a contradiction since $d_{\mathrm{ABW}}\not\equiv d_{\mathrm{KR}}$ from (5.3). ∎
@@ 6. Examples
In the following examples we let $\mu=\mathcal{N}(0,A)$ and $\nu=\mathcal{N}(0,B)$ be non-degenerate.
Example 6.1.
Suppose $A$ is diagonal. Then $L=A^{\frac{1}{2}}$ is also diagonal. For any $B$, we have $(L^{\top}M)_{tt}=L_{tt}M_{tt}>0$ for all $t$. By Corollary 4.6, the Knothe-Rosenblat coupling is $\mathcal{AW}_{2}$-optimal. This is a special case of [8, Theorem 2.9].
@@
(6.1) $A=\begin{bmatrix}1&2\\ 2&5\end{bmatrix}\quad\text{and}\quad B=\begin{bmatrix}1&-2\\ -2&5\end{bmatrix}.$
@@
Figure 2. Three interpolations $(A_{t})_{0\leq t\leq 1}$ between $A_{0}=A$ and $A_{1}=B$ as in (6.1). For visualization purposes we also include a uniform translation. The thin (grey) ellipses visualize McCann’s displacement interpolation. The thick dashed (red) ones show the Knothe-Rosenblatt interpolation where $T=L_{1}L_{0}^{-1}$. The thick solid (blue) ones show the interpolation with $T=L_{1}PL_{0}^{-1}$; in this case $A_{\frac{1}{2}}$ is degenerate.
Next we give an example where the optimal $\mathcal{AW}_{2}$-coupling is not unique.
@@
(6.2) $L=\begin{bmatrix}1&0\\ 1&1\end{bmatrix}\Rightarrow A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\quad\text{and}\quad M=\begin{bmatrix}1&0\\ -1&1\end{bmatrix}\Rightarrow B=\begin{bmatrix}1&-1\\ -1&2\end{bmatrix}.$
@@
(6.3) $X_{2}=X_{1}+\epsilon_{2},\quad Y_{2}=-Y_{1}+\epsilon_{2},$
where $\epsilon_{2}=\epsilon_{2}^{X}=\epsilon_{2}^{Y}\sim\mathcal{N}(0,1)$ is the common noise at time $2$. Since
@@
Since $\operatorname{diag}(L^{\top}M)$ has non-negative entries, the Knothe-Rosenblatt coupling is optimal. From (6.3), we see that $\mu$ is stochastically increasing but $\nu$ is stochastically decreasing, so $\mu$ and $\nu$ are not stochastically co-monotone (we refer the reader to [9, Definition 3.3] for definitions). In particular, stochastic co-monotonicity is not necessary for the Knothe-Rosenblatt coupling to be optimal, even if we restrict to Gauss-Markov processes.